Write a Pythagorean triplet whose one member is.  18

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

For any natural number m, we know that 2m, m2–1, m2+1 is a Pythagorean triplet.
2m=18
⇒m=18/2=9
m^2–1=9^2–1=81–1=80
m^2+1=9^2+1=81+1=82
∴ (18, 80, 82) is a Pythagorean triplet.

Thanks